point P(x,y) lies on the line joining A(3,4) and B(-5,-6) and also on the line joining C(0,-3), D(-2,1) find the cordinates of P

Grade:11

1 Answers

Arun
6 years ago
Dear Kashish
 
Line from A & B-
y -4 = (5/4) (x -3)
4y -16 = 5x - 15
4y -5x = 1
 
Now
Line from C & D-
y+3 = ( -2)(x-0)
2x + y = -3
Or
4y + 8x = -12
 
Now solve two equation
13 x = -13
x = -1
And also y = -1
 
Hence P = (-1, -1)
 
Regards
Arun(askIITians forum expert)